If P = 200 - 4Q, then the price elasticity of demand at P=80 is?
1
Expert's answer
2017-10-03T11:51:07-0400
If P = 200 - 4Q, P=80, then the price elasticity of demand is E=-b*(P/Q). Q = (80-200)/-4=30.
E= 4*(80/30)=10.67.